The area of the creative play space will only be used by the primary students. The total area of the yard is 800 m2 . Calculate the area of the yard that will be used by junior and intermediate students.



1
Expert's answer
2021-01-26T04:35:16-0500

The primary students will use "\dfrac{1}{4}" of the yard. This leaves "1-\dfrac{1}{4}=\dfrac{3}{4}" for the junior and intermediate students.

Each quarter represents "\dfrac{1}{4}\cdot800m^2=200m^2" of the yard.

Three of these sections would be "3\cdot200m^2=600m^2."

The area of the yard that will be used by junior and intermediate students is 600 m2.